OpenStudy (anonymous):

A salsa recipe uses green pepper, onion, and tomato in the extended ratio 2 : 5 : 9. How many cups of onion are needed to make 96 cups of salsa? A.12 B.5 C.30 D.54

OpenStudy (kropot72):

The proportion of onion is \[\frac{5}{2+5+9}=\frac{5}{16}\] Number of cups of onion needed to make 96 cups of salsa is \[96\times \frac{5}{96}=you\ can\ calculate\]